how do i do S=2(3.14)rh solve for h
Answer by mmm4444bot(95)

Hello There:
S = 2*(3.14)*r*h
The first thing we can do is simplify the expression on the right side of the equals sign because 2*(3.14) = 6.28
S = 6.28*r*h
To solve for h, we need to get rid of the 6.28*r on the right side because this will isolate h by itself (which is the goal of solving for a variable).
Divide both sides by 6.28*r.
S/(6.28*r) = (6.28*r)*h/(6.28*r)
h = S/(6.28*r)
~ Mark